A spring is stretched 40mm by a force of 15N. What is the work done by the force


1
Expert's answer
2022-03-28T11:02:47-0400

The work done by the force

"W=\frac{kx^2}{2}=\frac{Fx}{2}\\
=\frac{15\:\rm N*0.04\;\rm m}{2}=0.3\:\rm J"